The function f: [ 2,4] to R , defined by f(x)= 3/x is uniformly continuous on its domain.


True or false with full explanation

1
Expert's answer
2021-11-18T06:54:20-0500

Define "f(x)=\frac{3}{x} \forall x \in[2,4]."

Then f is continuous and bounded function on [2,4].

For any "\delta>0" , we can find "m \in \mathbb{N}" such that "\frac{1}{n}<\delta" for every "n \geq m"

Let "x_{1}=\frac{3}{m}, x_{2}=\frac{3}{2 m}" , so that "x_{1}, x_{2} \in[2,4]"

Consider "\left|x_{2}-x_{1}\right|=\left|\frac{3}{2 m}-\frac{3}{m}\right|=\frac{3}{2 m}>\frac{\delta}{2}>\delta"

and "\left|f\left(x_{2}\right)-f\left(x_{1}\right)\right|=|\frac{2 m}{3}-\frac{m}{3}|=\frac{m}{3}"

Therefore if we choose "\varepsilon>0" such that "\varepsilon>m"

then we get "\left|x_{2}-x_{1}\right|>\delta" and "\left|f\left(x_{2}\right)-f\left(x_{1}\right)\right|<\varepsilon"

Hence f is uniformly continuous on [2,4].